OpenStudy (anonymous):

Jason had $32. He spent all the money on buying 4 CDs for $x each and 2 magazines for $y each. If Jason would have bought 5 CDs and 2 magazines, he would have run short of $4. A student concluded that the price of each CD is $5 and the price of each magazine is $6. Which statement best justifies whether the student's conclusion is correct or incorrect?

OpenStudy (anonymous):

ANSWER CHOICES The student's conclusion is incorrect because the solution to the system of equations 4x + 2y = 4 and 5x + 2y = 36 is (9, 4). The student's conclusion is incorrect because the solution to the system of equations 4x + 2y = 32 and 5x + 2y = 36 is (4, 8). The student's conclusion is correct because the solution to the system of equations 4x + 2y = 4 and 5x + 2y = 36 is (5, 6). The student's conclusion is correct because the solution to the system of equations 4x + 2y = 32 and 5x + 2y = 36 is (5, 6).

OpenStudy (anonymous):

@ranga